You are the manager of a monopolistically competitive firm. The present demand curve you face is P = 100 – 4Q. Your cost function is C(Q) = 50 + 8.5Q2.

                       i.        What level of output should you choose to maximize profits?

                      ii.        What price should you charge?

                    iii.        What will happen in your market in the long run? Explain.


1
Expert's answer
2021-11-01T19:22:58-0400

A monopolistically competitive firm maximizes profits by equating its MR with MC:

100 - 8Q = 17Q

25Q = 100

Q = 100 / 25 = 4 (Profit-maximizing output)

P = 100 - 4Q = 100 - (4 x 4) = 100 - 16 = 84 (Profit-maximizing price)

In the long run, these equilibrium values will not be maintained. A monopolistically competitive firm has to differentiate its product from that of the other firms and in an attempt to do so, has to resort to different differentiating techniques. Over time, this will keep increasing the firm's total cost. As total cost increases, ceteris paribs its profit will start decreasing. So, in the long run, the firm will earn lower profit (or loss, deprending on its cost & pricing relationship) and may choose to exit the market.